OpenStudy (anonymous):

The length of a rectangle is 1 M less than twice its width, and the area of the rectangle is 55 M^2 . Find the dimensions of the rectangle.

terenzreignz (terenzreignz):

First step to such word problems is represent We let w = width of the rectangle, what then is the length?

terenzreignz (terenzreignz):

The length is 2w - 1, since the length is 1 M less than twice its width

terenzreignz (terenzreignz):

Area = 55 = w(2w - 1) 0 = 2w^2 - w - 55 w = [1 + sqrt (1 + 440)]/4 w = (1 + 21)/4 w = (22/4) = 5.5 w = 5.5 is the width 2w - 1 = 10 is the length